OpenStudy (anonymous):

As a sound source approaches someone, the frequency of the sound wave increases. decreases. decreases, then increases. remains constant.

OpenStudy (anonymous):

remains constant as long as either of two involves no movement..

OpenStudy (anonymous):

So, it is not A?

OpenStudy (anonymous):

no, it is d because it hasn't said the sound SOURCE involved movement: it is fixed and therefore no change of frequency would expect.

OpenStudy (anonymous):

It is like a wave spreading out with a constant frequency.

OpenStudy (anonymous):

Remember that frequency and wavelength are inversely proportional to each other. It all depends on the situation here. Remember that: v = f * λ where... f = frequency λ = wavelength ....................................................................

OpenStudy (anonymous):

@aryandecoolest it is not always the case..

OpenStudy (anonymous):

yeah i know!!.....Based on Doppler effect. when the observer is moving relative to the wave source, the observer experiences the Doppler Effect. The motion of the source causes a change in the sound's wavelength. Wavelength is inversely proportional to frequency. The apparent frequency is given by the equation:f = (v-vr)/(v-vs ) * f0where f is the apparent frequency, v is the wave's velocity, vs is the velocity of the source relative to the medium, and is vr the velocity of the observer relative to the medium.Imagine a police siren as the car moves toward, then away from you. We've all heard that, right? The siren's frequency increases, passes you, then decreases. It's the relative motion between you and the car that causes your ears to discern a difference in pitch....